Mar 26, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. BCAL pedestals: Tegan (UofR) will summarize the determination of pedestals in a note with plots in preparation for the collaboration meeting.
  2. Wish-list of action items: discuss attached document (Media:BCAL_Action_Items-2015.pdf) and add chosen items to Wiki page. Zisis will update the list
  3. Will: Create histograms to determine number of pi0s and pi0 width for online monitoring.
  4. Work with Hovanes to activate VME LED pulser setup for FCAL (Elton)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Spring Run
  4. FCAL Update
  5. BCAL Update
  6. Reconstruction and Simulation
  7. Calibration
  8. Any other business


Attendees: Elton, Will, Mark, Adesh (JLab); Matt (IU); Sean (NW); Tegan, Zisis (UofR); Christina, George (Athens)

  1. Announcements
    • Power outage of accelerator site yesterday due to VA Power transformer malfunction.
    • Power lost late afternoon, but most systems were back up by early evening. LCW systems were restored by 9 pm last night.
    • Recover of systems began this morning and will continue through the day. Accelerator hopes to start restoring beam later this evening.
    • Hall D powered down as expected, with only a few glitches. The ODH and PSS systems should have been on backup power but were not.
    • The solenoid weathered the power outage gracefully. This morning they were exercising the system with low current
    • Other systems are being restored. BCAL has been powered back. Elton and Adesh will get the FCAL back this afternoon
  2. Action Items
    • Most addressed as some level.
    • Zisis' list will be moved to ongoing activities.
  3. Spring Run
    • To date, we have not received clean beam. Accelerator has delivered beam to the tagger dump, but radiation levels were between 3-20 times comparable levels last fall.
    • The accelerator has commissioned the RF separation system, so they should be able to deliver beam to Hall A and D simultaneously when they come back.
  4. FCAL Update
    • Manuel has made mods to the HV GUI displaying the entire array. This is a great improvement.
    • Some discussion regarding viewpoint for display. Nominal x to the left is considered, but Elton noted that hdview2 has x to the right. So they might make it compatible.
    • Adesh has attempted gain balancing with electrons. Showers are selected based on the track projection to a particular block and uses the summed energy of 25 surrounding blocks.
    • Adjustments based on the electrons has made marginal improvements in resolution, similar to those obtained by smoothing the LED pulser response.
    • Online monitoring plot shows 2g invariant mass and could be used to estimate number of pi0s online. Sean: Which particular plot should be used, as there are several? Matt: Look in the cluster space and select the 2g invariant mass.
    • Adesh has also worked with Sasha on investigating trigger thresholds with cosmic-ray data. They note that a threshold of 60 counts significantly reduces the FCAL rate. Waiting on beam for further studies.
  5. BCAL Update
    1. See Log 3328508 and 3324355. (Will) Calibration update.
      • Plots of gain calibration constants. Moved input peak (mainly layer 4) to be more representative (i.e. central).
      • Mark: (Will add a log entry later). Investigated an alternative technique for gain calibration: Plot 2g invariant mass whenever a cell contributes to one of the showers and look for systematics. The plots show that very few pi0s are showing up in layer 4, which is why pi0 width minimization does not seem to affect these constants.
      • Christina: Is it possible that the showers are not reaching the back of the BCAL? Are there higher energy pi0s (e.g. 3 GeV?). Will: Not much data for Eg>1.5 GeV.
      • Zisis: Simulations of showers exist and he will look at them for guidance.
    2. Single-ended hits (Elton)
      • See Log 3330427
      • Inclusion of double ended hits into cluster adds very little
      • Inclusion of single-ended hits adds of order 3%
      • Mark: Changes to the clusterizer to include single-ended hits should involve a systematic study of many aspects of cluster finder and include diagnostics, validation and testing.
      • Aspects that should be studied include: neutral vs charged track showers. a) Note that clusterizer was optimized for photons, but Elton is looking at showers matched to charged tracks. b) luminosity effects. At the low luminosity of last fall, overlap of clusters is rare, but cluster optimization should consider situations at high rates where it may be more difficult to separate out showers.
      • These issues should be taken over by the BCAL reconstruction meeting.
  6. Reconstruction and Simulation
    • (Tegan): Clustering diagnostics to investigate shower curvature observed in MC by andrei are underway. Bugs are being found and fixed.
  7. Calibration
    • (Elton) Gain calibration constants for BCAL in ccdb are based on the 1.2 V overbias setting from last fall. At the moment the biases are set to 1.4 V overbias. Early in the run, we will take some voltage scans to determine the optimum setting for the overbias. If they remain at 1.4 V, an initial set of calibration constants for the spring run should be added to ccdb by scaling the current constants by 1.4/1.2.
  8. Any other business